OpenStudy (anonymous):

PLEASE HELP, WILL REWARD! Does the infinite geometric series diverge or converge? Explain. 2+6+18+54+.....

hartnn (hartnn):

can you find the common ratio "r" of this geometric series ? if r >1 , the geometric series diverges. if r < 1, the geometric series converges

OpenStudy (anonymous):

Is it r<1?

OpenStudy (anonymous):

Do you know if it has a sum or doesn't have a sum?

hartnn (hartnn):

common ratio is the ratio between the terms, like 2nd term/ first term = 3rd term/2nd term here r= 6/2 = 18/6 =.... = 3 and since r = 3 r>1

hartnn (hartnn):

so the sum diverges and the series does not have sum